A major farming company is planning to purchase new tractors. It wants to borrow $650 million by issuing bonds. The bonds are for a 12-year period with simple interest computed quarterly at a rate of 10 percent per year. Interest is to be paid each quarter to bondholders. How much will the airline have to pay it. quarterly interest? How much interest will it pay in the 12-year period?


1
Expert's answer
2021-10-11T16:46:52-0400
"650\times (1+\frac{1}{2.5^4})^{40}=1786.6"

"1786.6-650=1136.6"